Can somebody help with this?
Experiment 1 - Determination of Local Free-Fall Acceleration Name Lab Day Lab Time wiiO Partners Du Omoled Lab Instructor piup e Procedure A Initial distance of fall rn ± 0.002 m Proportional error in distance 부 Times of Fall (in seconds): Drop l Drop 5 Average time of fall Sample standard deviation in time-of-fall data Standard error in the mean for time of fall Table 1-1 value for η Uncertainty in the time of fall Proportional error in time Calculated value of g Proportional error in g Uncertainty in g for a level of confidence of_ - - m/s m/s 1-5
Answer #1

(i) The average time of fall which will be given by -

\bar{t}avg = (t1 + t2 + t3 + t4 + t5) / (N)

\bar{t}avg = [(0.6047 s) + (0.6045 s) + (0.6043 s) + (0.6046 s) + (0.5989 s)] / (5)

\bar{t}avg = [(3.017 s) / (5)]

\bar{t}avg = 0.6034 s

(ii) The sample standard deviation in time-of-fall data which will be given by -

\sigmat = \sqrt{}(1/N) \sum_{i=1}^{N} (ti - \bar{t}avg)2

where, (t1 - \bar{t}avg)2 = [(0.6047 s) - (0.6034 s)]2 = 1.69 x 10-6 s2

(t2 - \bar{t}avg)2 = [(0.6045 s) - (0.6034 s)]2 = 1.21 x 10-6 s2

(t3 - \bar{t}avg)2 = [(0.6043 s) - (0.6034 s)]2 = 8.1 x 10-7 s2

(t4 - \bar{t}avg)2 = [(0.6046 s) - (0.6034 s)]2 = 1.44 x 10-6 s2

(t5 - \bar{t}avg)2 = [(0.5989 s) - (0.6034 s)]2 = 2.025 x 10-5 s2

then, we get

\sigmat = \sqrt{}(1/5) [(1.69 x 10-6 s2) + (1.21 x 10-6 s2) + (8.1 x 10-7 s2) + (1.44 x 10-6 s2) + (2.025 x 10-5 s2)]

\sigmat = \sqrt{}(1/5) (2.54 x 10-5 s2)

\sigmat = \sqrt{}5.08 x 10-6 s2

\sigmat = 0.00225 s

(iii) The standard error in mean for time-of-fall which will be given by -

\sigmaavg = \sigmat / \sqrt{}N

\sigmaavg = (0.00225 s) / \sqrt{}5

\sigmaavg = [(0.00225 s) / (2.236)]

\sigmaavg = 0.001 s

(v) The uncertainty in time-of-fall which will be given as -

Uncertainty = \bar{t}avg\pm\sigmat

Uncertainty = (0.6034 s) \pm(0.00225 s)

(vi) The proportional error in time which will be given by -

Error = \sigmat / \bar{t}avg

Error = [(0.00225 s) / (0.6034 s)]

Error = 0.00372